Stefan Ingves, governor of the Swedish National Bank (Riksbanken), reportedly was paid 14 million SEK in advance to write his book How we boosted inflation. The book took three years to write. In the time he spent writing, Stefan could have been paid to work as a professor and give speeches as an economic commentator in TV shows. Given his knowledge and experience, assume that he could have earned 8 million SEK per year (paid at the end of the year) working and commenting instead of writing the book. Assume his cost of capital is 6.0% per year.

a. What is the NPV of agreeing to write the book (ignoring any royalty payments)?

The NPV is $ _________million. (Round to three decimals)

b. Assume that, once the book is finished, it is expected to generate royalties of $5 million in the first year (paid at the end of the year) and these royalties are expected to decrease at a rate of 30% per year in perpetuity. What is the NPV of the book with the royalty payments?

The NPV of the book with the royalty payments is ________$ million. (Round to three decimals, DO NOT ROUND THE RESULT FROM (a))

c. How many IRRs are there in part (a)?

The number of IRRs is _____.

Does the IRR rule give the right answer in this case?

Fill in "yes" or "no"

d. How many IRRs are there in part (b)?

The number of IRRs is _________.

Does the IRR rule give the right answer in this case?

Fill in "yes" or "no" .
Answer #1
a A B C = A x B
Year CF DF at 6% Present Value
0 14 1 14
1 -8 0.943396 -7.547169811
2 -8 0.889996 -7.11997152
3 -8 0.839619 -6.716954264
NPV                 (7.384) Million
$8 Million cash outflow is opportunity foregone for three years working as professor
So NPV of agreeing to write books will have negative NPV of 7.384 million
b Since royalty is started when book is finished in 3 years i.e. its started in 4th year onwards
with declining perpetuity
Present value of perpetuity royalty at year 3 = amount of royalty / (R - d)
5/(0.06-(-0.3))
13.88889 million
PV of royalty = 13.8889/1.06^3
11.661 million
NPV with royalty = -7.384+11.661 = 4.277 million
c Year CF
0 14
1 -8
2 -8
3 -8
IRR = 32.68% There is only one Irr of 32.68% that is higher than discount rate
IRR(values 0 to 3) But NPV is negative so IRR gives wrong decision, NO
d In this case there is 2 irr since royalty amount turning the cash flow sign from + to - and again +
No, IRR rule does not provide correct answer since two IRR is there.